# Mimic In-Context Learning for Multimodal Tasks

### v1

@@ 1 Introduction
Due to synergistic effects, when implementing ICL, it is hard for LMMs to capture the general mapping from input-output pairs of complex multimodal ICDs as LLMs. Instead, [23] shows that LMMs tend to rely on the distribution of ICDs to narrow the prediction space, e.g., in Visual Question Answering (VQA), a LMM might recognize the ICD answer format and respond an answer in the same format, rather than learning the correct function as in language QA. Consequently, the ICL performance in LMMs is more sensitive than in LLMs to ICD configurations [28, 58] and finding optimal ICD configurations in LMM is still an open question [7]. A straightforward approach to mitigate the high sensitivity issue is to use more ICDs to help LMMs recognize stable patterns for improved predictions. However, image inputs require more tokens than text, and increasing the number of ICDs significantly raises computational demands. Moreover, current LMMs, like open-source 8B models, typically support only up to 32-shot ICDs [18], rendering this approach impractical at larger scales.
@@ 3.1 Mathematic Analyses
$=\frac{Z_{2}}{Z_{1}+Z_{2}}\textrm{softmax}(\bm{q}\bm{K}^{\top})\bm{V}+\frac{Z_{1}}{Z_{1}+Z_{2}}\exp(\bm{q}\bm{K}_{D}^{\top})\bm{V}_{D}$
$=(1-\mu)\textrm{SA}(\bm{q},\bm{K},\bm{V})+\mu\textrm{SA}(\bm{q},\bm{K}_{D},\bm{V}_{D})$ (2)
@@ 3.2 Mimicking ICD Affected Terms
To approximate $Z_{1}(\bm{q},\bm{K}_{D})$, we note that it is a positive scalar dependent solely on the current query token $\bm{q}$ and the ICD keys $\bm{K}_{D}$. Therefore, we use a simple mapping: a trainable linear layer $f(\cdot):\mathbb{R}^{d_{h}}\to\mathbb{R}$ to approximate $\log Z_{1}$.ntion difference term $\textrm{SA}(\bm{q},\bm{K}_{D},\bm{V}_{D})-\textrm{SA}(\bm{q},\bm{K},\bm{V})$, as Fig. 2(a) shows, we insert a learnable vector $\bm{v}\in\mathbb{R}^{d_{h}}$ in each attention head to capture the general shift effect for this head. Then, the output of Eq. 2 in MimIC attention head is computed as $\operatorname{SA}(\bm{q},\bm{K},\bm{V})+\tilde{\mu}(\bm{q},\bm{K})\bm{v}$, where $\tilde{\mu}(\bm{q},\bm{K})=\tilde{Z_{1}}(\bm{q})/(\tilde{Z_{1}}(\bm{q})+Z_{2}(\bm{q},\bm{K}))$ and $\tilde{Z_{1}}(\bm{q})=\exp(f(\bm{q}))$. After obtaining the outputs from all MimIC attention heads, they are concatenated, flattened, and passed through the matrix $\bm{W}_{o}\in\mathbb{R}^{d\times d}$ and FFN layer.
@@ 5 Conclusion
Motivated by the insight that in-context demonstrations function as shift vectors applied to the hidden states of query tokens, we propose Mimic In-Context Learning (MimIC) to learn this effect in Large Multimodal Models (LMMs). MimIC operates by inserting distinct shift vectors into different heads of attention layers, employing a linear layer to generate query-dependent magnitudes for these shift vectors, and leveraging a layer-wise alignment loss to align with ICL. Empirical evaluations across three diverse tasks using two LMMs demonstrate MimIC achieves competitive few-shot in-context learning performance with significantly reduced inference latency compared to traditional ICL methods, requires fewer training samples relative to LIVE, consistently surpasses other shift vector-based approaches, utilizes fewer parameters than LoRA yet yields superior results, and substantially reduces hallucinations compared to existing techniques.
@@ 6 Acknowledgment
This work is supported by the National Science Foundation of China (62206048), the Natural Science Foundation of Jiangsu Province (BK20220819), and the Fundamental Research Funds for the Central Universities (2242024k30035). This research work is also supported by the Big Data Computing Center of Southeast University

### v2

@@ 1 Introduction
Due to synergistic effects, when implementing ICL, it is hard for LMMs to capture the general mapping from input-output pairs of complex multimodal ICDs as LLMs. Instead, [23] shows that LMMs tend to rely on the distribution of ICDs to narrow the prediction space, e.g., in Visual Question Answering (VQA), a LMM might recognize the ICD answer format and respond an answer in the same format, rather than learning the correct function as in language QA. Consequently, the ICL performance in LMMs is more sensitive than in LLMs to ICD configurations [28, 58] and finding optimal ICD configurations in LMM is still an open question [7]. A straightforward approach to mitigate the high sensitivity issue is to use more ICDs to help LMMs recognize stable patterns for improved predictions. However, image inputs require more tokens than text, and increasing the number of ICDs significantly raises computational demands. Moreover, current LMMs, like open-source 9B models, typically support only up to 32-shot ICDs [18], rendering this approach impractical at larger scales.
@@ 3.1 Mathematic Analyses
$=\frac{Z_{2}}{Z_{1}+Z_{2}}\textrm{softmax}(\bm{q}\bm{K}^{\top})\bm{V}+\frac{Z_{1}}{Z_{1}+Z_{2}}\textrm{softmax}(\bm{q}\bm{K}_{D}^{\top})\bm{V}_{D}$
$=(1-\mu)\textrm{SA}(\bm{q},\bm{K},\bm{V})+\mu\textrm{SA}(\bm{q},\bm{K}_{D},\bm{V}_{D})$
$=\underbrace{\textrm{SA}(\bm{q},\bm{K},\bm{V})}_{\textrm{standardattention}}+\underbrace{\mu(\textrm{SA}(\bm{q},\bm{K}_{D},\bm{V}_{D}))-\textrm{SA}(\bm{q},\bm{K},\bm{V})}_{\textrm{shift vector}}$ (2)
@@ 3.2 Mimicking ICD Affected Terms
To approximate $Z_{1}(\bm{q},\bm{K}_{D})$, we note that it is a positive scalar dependent solely on the current query token $\bm{q}$ and the ICD keys $\bm{K}_{D}$. Therefore, we use a simple mapping: a trainable linear layer $f(\cdot):\mathbb{R}^{d_{h}}\to\mathbb{R}$ to approximate $\log Z_{1}$. For the attention difference term $\textrm{SA}(\bm{q},\bm{K}_{D},\bm{V}_{D})-\textrm{SA}(\bm{q},\bm{K},\bm{V})$, as Fig. 2(a) shows, we insert a learnable vector $\bm{v}\in\mathbb{R}^{d_{h}}$ in each attention head to capture the general shift effect for this head. Then, the output of Eq. 2 in MimIC attention head is computed as $\operatorname{SA}(\bm{q},\bm{K},\bm{V})+\tilde{\mu}(\bm{q},\bm{K})\bm{v}$, where $\tilde{\mu}(\bm{q},\bm{K})=\tilde{Z_{1}}(\bm{q})/(\tilde{Z_{1}}(\bm{q})+Z_{2}(\bm{q},\bm{K}))$ and $\tilde{Z_{1}}(\bm{q})=\exp(f(\bm{q}))$. After obtaining the outputs from all MimIC attention heads, they are concatenated, flattened, and passed through the matrix $\bm{W}_{o}\in\mathbb{R}^{d\times d}$ and FFN layer.
@@ 5 Conclusion
Motivated by the insight that in-context demonstrations function as shift vectors applied to the hidden states of query tokens, we propose Mimic In-Context Learning (MimIC) to learn this effect in Large Multimodal Models (LMMs). MimIC operates by inserting distinct shift vectors into different heads of attention layers, employing a linear layer to generate query-dependent magnitudes for these shift vectors, and leveraging a layer-wise alignment loss to align with in-context learning (ICL). Empirical evaluations across three diverse tasks using two LMMs demonstrate MimIC achieves competitive few-shot ICL performance with significantly reduced inference latency compared to traditional ICL methods, requires fewer training samples relative to LIVE, consistently surpasses other shift vector-based approaches, utilizes fewer parameters than LoRA yet yields superior results, and substantially reduces hallucinations compared to existing techniques.
@@ 6 Acknowledgment
This work is supported by the National Science Foundation of China (62206048), the Natural Science Foundation of Jiangsu Province (BK20220819), and the Fundamental Research Funds for the Central Universities (2242024k30035). This research work is also supported by the Big Data Computing Center of Southeast University.